OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (anonymous):

@hartnn can you help me with this one?

OpenStudy (yamyam70):

is it sin squared x ? sin^2 x

OpenStudy (anonymous):

Yes sorry about that, should be: 2 sin^2x = sin x

OpenStudy (yamyam70):

okay first we equate it to zero. then we factor the Left hand side.

OpenStudy (anonymous):

Okay, can you do everything at once so i can see each step by step? and then if i dont understand it i can ask for an explanation? Thanks

OpenStudy (yamyam70):

I just can't give answers, but don't worry I'll guide you :)

OpenStudy (anonymous):

Well this is only a practice problem so its not like it counts towards anything, i have several other problems i can do on my own. It would really help me if it was all laid out how to do it and then i can do some on my own understanding it all.

OpenStudy (yamyam70):

Have you done any work so far?

OpenStudy (anonymous):

Not on this problem.

OpenStudy (yamyam70):

I mean with the given practice problem.

OpenStudy (yamyam70):

okay first things first , we must equate it to 0

OpenStudy (yamyam70):

and then factor the left hand side.

OpenStudy (yamyam70):

if we equate 2 sin^2 x = sin x , it will be 2 sin ^2 x - sin x = 0

OpenStudy (yamyam70):

do you follow?

OpenStudy (anonymous):

So far, yes

OpenStudy (yamyam70):

can you factor the left hand side?

OpenStudy (yamyam70):

TIP : think of sin^2 x as = x^2

OpenStudy (anonymous):

Im not sure how to factor this one because of it saying sin and sin x?

OpenStudy (yamyam70):

okay for you not to be confused , think of

OpenStudy (yamyam70):

\[\sin^2x = x^2\]

OpenStudy (yamyam70):

sin x = x

OpenStudy (anonymous):

Okay, but then whats there to factor?

OpenStudy (anonymous):

Im sorry, see this is why it would be easier if you did it all step by step and then i ask questions

OpenStudy (yamyam70):

we can factor this out by taking out what is common, first substitute

OpenStudy (yamyam70):

\[2 \sin^2 x - \sin x \]

OpenStudy (yamyam70):

to ,

OpenStudy (yamyam70):

\[2 x^2 - x \]

OpenStudy (yamyam70):

then factor out what is common

OpenStudy (anonymous):

the x's?

OpenStudy (yamyam70):

correcT :)

OpenStudy (anonymous):

Okay so whats next? sorry im kind of in a hurry to finish this problem to do the rest

OpenStudy (yamyam70):

okay I'm just going to show you

OpenStudy (yamyam70):

x ( 2x -1 ) = 0 substitute x with sin x ( x = sin x , x^2 = sin^2 x ) sin x ( 2 sin x -1 ) = 0 sin x = 0 , ..................... 2 sin x -1 = 0 ........... 2 sin x = 1 ,....... sin x = 1/2 Check by substition.

OpenStudy (yamyam70):

note in checking the we must equal both sides. so , first 2 sin^2 x = sin x 2 ( 0 )^2 = 0 , 2 (1/2)^2 = 1/2, If both are equal its therefore the solution .

OpenStudy (anonymous):

So the answer is 1/2?

OpenStudy (yamyam70):

is it only 1/2 ?

OpenStudy (anonymous):

wait, 0, 1, and 1/2 right?

OpenStudy (yamyam70):

0 and 1/2 is the solution set.

OpenStudy (anonymous):

Ohhhh i see it now

OpenStudy (yamyam70):

Be sure to learn it and practice :) I got to go bye :)

OpenStudy (anonymous):

@yamyam70 did you solve this problem? sin2x + sin x = 0

OpenStudy (yamyam70):

there is no such thing :)

OpenStudy (yamyam70):

should be , sin ^2 x + sin x = 0

OpenStudy (anonymous):

This was the problem i needed solved, is this the problem we solved? Find all solutions to the equation. sin^2x + sin x = 0

OpenStudy (anonymous):

Because earlier i replied incorrectly and said: 2 sin^2x = sin x

OpenStudy (yamyam70):

its the same thing, much easier, you can do it :)

OpenStudy (anonymous):

No, im asking which one you solved.

OpenStudy (yamyam70):

Please go back to what I have solved earlier.

OpenStudy (anonymous):

You solved this problem 2 sin^2 x = sin x , i needed this problem solved sin^z x = sin x

OpenStudy (anonymous):

Sin^2 x = sin x*****

OpenStudy (anonymous):

I had a typo and now we did the wrong problem :(

OpenStudy (yamyam70):

If you look closely. There is no big difference between 2 sin^2 - sin x = 0 and sin^2 - sin x = 0

OpenStudy (yamyam70):

now you have your example , please do sin^x - sin x = 0 :)

OpenStudy (yamyam70):

*sin^2x - sin x = 0

OpenStudy (anonymous):

Okay i will, but can you tell me which one of these answers is correct for this problem: Find all solutions in the interval [0, 2π). 2 sin2x = sin x A. x = pi/3 , 2pi/3 B. x = pi/2 , 3pi/2 , pi/3 , 2pi/3 C. x = 0 , pi , pi/6 , 5pi/6 D. x = pi/6 , 5pi/6

OpenStudy (anonymous):

We solved 2 sin^2x = sin x together but we didnt do it in the interval [0, 2π).